Instrument Log 

RSS                 C1                  CF Lamont, OK

Date

Technician

Comments

10/1/14

CAM

The dry air system compressor motor failed on 9/30.  Removed the failed motor and replaced with a spare rebuilt motor.  Removed smaller compressor that had been temporarily plumbed in yesterday and placed the original compressor back in service at 2000 GMT.

9/30/14

CAM

Technician found during daily rounds that the compressor motor for the dry air system had failed.  Temporarily connected a smaller compressor to the dry air system until the large compressor can be repaired.  RTS at 2115 GMT.  Note: RSS is not currently operational.

3/13/14

JEM

Mentor indicates the RSS has been down since a recent power outage.  Observed the RSS still rotating and shading properly.  Opened the instrument enclosure and there was a strong odor of burnt electronics.  Further inspection revealed that the RSS control board was heavily damaged beyond repair.  Notified the mentor what was found and told SDS staff to disable collections.  RSS will remain down until a replacement board is received.  Did not RTS.

1/9/14

JEM

The portable calibration and vertical HgCd calibration need to be performed.  Performed the vertical HgCd calibration between 2015 GMT and 2022 GMT.  Performed the portable calibration between 2024 GMT and 2108 GMT.  RTS @2110 GMT.

12/26/13

CAM

Shadow band was running, but not shading properly.  Adjusted shadow band at 1658 GMT.  RTS status O.K.

12/11/13

JEM

Bi-weekly portable calibration needs to be performed.  Performed calibration between 1724 GMT and 1804 GMT.  RTS @1804 GMT.

11/27/13

JEM

The RSS portable calibration is scheduled.  Connected the calibrator control cable at 2003 GMT.  Mounted the calibrator and ran the calibrator from 2005 GMT to 2030 GMT.  Removed the calibrator and disconnected the calibrator control cable at 2032 GMT.  RTS @2035 GMT.

9/25/13

JEM

The RSS has not been transferring any data since 9/17/13.  Found the RSS operating normally.  Checked the computer and the log file had not updated since 9/17 and the most recent data file was dated 9/17.  Checked that xtty was running and it was.  Killed the xtty process and restarted the xtty process @1541 GMT.  Checked the log file and data file and both were increasing in size.  RTS @1600 GMT.

7/12/13

KT

The shadow band was not rotating. At 1325 GMT power cycled instrument. At 1329 GMT the shadow band began normal operations. RTS at 1340 GMT.

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Powered UPS on at 1945 GMT.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

1/14/13

CAM

After reinstalling the SWS, noticed the dry air system was not providing air.  T/S found the copper air line between the compressor and the building was plugged with ice.  Disconnected copper line at building entrance, thawed the ice, and cleared water from line.  Reconnected line and confirmed air flow.  RTS at 2030 GMT.   SAS-He, RSS and SORTI were also affected by lack of air.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Power was restored at 2115 GMT after electrical upgrades were complete. 

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  RSS auto-started status O.K.

8/27/12

PAD

Last data plots shown for RSS is on 8/25/12 at 1040 GMT.  At 1438 GMT, found instrument without power.  Troubleshooting found UPS located in Optical trailer had tripped the breaker.  Reset breaker and power was restored to instrument.  Instrument powered up and was shading correctly soon after.  RTS at 1508 GMT.

6/21/12

PAD

Shutter control board has been repaired and returned from mentor.  Mentor requests board installed and instrument powered up.  Installed board at 1837 GMT.  Powered up instrument at 1919 GMT and checked replacement resistor on board for heat.  Resistor was cool to touch five minutes after powering up instrument.  Shadowband became active within three minutes after power was restored.  Unable to check shading due to sky conditions.  Notified mentor.  Board model YES RSS Control Board, SN YPN021100-1 Rev. 2.  RTS at 1927 GMT.

5/15/2012

PAD

Previous troubleshooting found a burnt resistor on RSS control board.  Mentor requests board removed and sent to him for repairs.  Shut off ion pump and main breaker to instrument at 1749 GMT.  Removed control board and took to shipping dept.  Restored power to instrument enclosure and powered up ion pump.  Board model YES RSS Control Board, SN YPN021100-1 Rev. 2.  RTS at 1829 GMT.

5/14/12

PAD

RSS has an issue of tripping power/breaker switch when powering up system.  Something got hot on recent restart and small amount of smoke was seen inside enclosure on power up but was unable at the time to pinpoint source.  Received troubleshooting instructions from mentor.  Checked voltages on power connectors for control and comms boards.  Found burnt resistor on control board.  Recorded voltages with both boards disconnected, comms board connected, and control board connected.  Took pictures of burnt resistor on control board.  Sent information to mentor.  Waiting on reply at this time.  RTS at 1830 GMT.

5/2/12

PAD

Instrument has been tripping breaker and shutting down.  Sometimes instrument will trip breaker when restarting instrument.  Mentor requests instrument shut down until issue can be investigated.  Shut down instrument at 1840 GMT.  Troubleshooting with mentor at this time.  RTS at 1855 GMT.

4/30/12

PAD/CAM

Shadowband was not operating at 1631 GMT.  Opened enclosure and observed instrument had power.  Shut off/ turned on power/breaker switch.  About 5 seconds after powering instrument up, I smelled then saw smoke coming from area behind CCD chamber and in front of RSS controller board.  I shut off power switch at that time.  At 1845 we inspected boards but did not see any indication of what was getting hot.  Disconnected power to Comms board and powered up.  Breaker did not trip.  Reconnected comms board and powered up instrument.  Breaker tripped the first time then instrument stayed powered up on retry.  No indication of anything getting hot at this time.  Notified instrument mentor.  RTS at 1900 GMT.

4/26/12

PAD/KT

New flexible ducts were ordered to be installed between external air conditioning unit and RSS enclosure.  Installed ducts and powered up AC unit.  RTS at 1827 GMT.

4/11/12

PAD

During PM rounds observed shadowband is not active.  At 1551 GMT, found instrument power switch tripped.  Reset switch.  Observed shadowband active at 1556 GMT.

4/10/12

PAD

At 1612 during PM rounds, observed shadowband was staying in home position.  At 1811 GMT, found shadowband was at home and not shading but had power.  Rebooted computer and attempted to power cycle instrument.  Power switch for instrument that also serves as a breaker would trip within a second after instrument was powered up.  Instrument failed to power up after several attempts.  Notified mentor.  At 1855 GMT, tried to power up instrument again and instrument stayed powered up after second attempt.  Shadowband was active shortly after start up.  Unable to verify shading due to overcast conditions.  RTS at 1901 GMT.

10/12/11

PAD

Last SDS collection was on 10/11/11 at 2100 GMT.  Shadowband is not active.  Powercycled instrument and shadowband is now shading correctly.  RTS at 1436 GMT.

10/11/11

PAD

Last SDS collection was on 10/9/11 at 2000 GMT.  Shadowband is not active.  Powercycled RSS unit at 1453 GMT.  After troubleshooting unrelated issue with another instrument, observed at 2010 GMT RSS shadowband was still not operational.  Rebooted computer and shadowband is now shading correctly.  RTS at 2025 GMT.

9/26/11

PAD

Purge Gas Generator located in Optical trailer has been repaired.  Unit needs to be RTS.  Installed Whatman FT-IR Model 75-45-12VDC, SN 7545-0236 Purge Gas Generator and connected inlet line.  Started unit and opened inlet valve to allow air to flow through system.  Vendor recommends unit allow to run 24 hours before connecting outlet line.  Will connect outlet line tomorrow.  Instruments affected are SWS, RSS, and SORTI.  RTS at 2120 GMT.

9/1/11

SPD

RSS shadow band was dead.  Found that the instrument power switch/circuit breaker had tripped to the off position. Turned switch to on position and it tripped again. Once more turned switch to on position and it stayed on. RSS powered up and shadow band shaded the diffuser.  RTS at 1445 GMT.

8/31/11

SPD

RSS shadow band was dead.  Found that the instrument power switch/circuit breaker had tripped to the off position. Turned switch to on position and it tripped again. Once more turned switch to on position and it stayed on. RSS powered up and shadow band shaded the diffuser.  RTS at 1624 GMT.

8/23/11

PAD

Sporadic comms issue between computer and instrument has been going on since 8/1/11.  Last SDS collection was on 8/21/11 at 0900 GMT.  Instrument was powercycled on 8/22/11 but it did not establish comms with computer.  Bit driver transmit light is on solid at instrument.  Unable to connect to comm/control board with laptop using ProComm.  Tested comms with loopback and found no problems.  Troubleshooting with mentor and found RS232 analyzer shows #3 Receive light is light on analyzer when connected directly to the board.  Powercycled instrument and connected to RSS computer.  At 2030 GMT, observed instrument was active and shading correctly.  Mentor is sending replacement RS232 chips for control/comm board.  Troubleshooting at this time.

8/22/11

SPD

RSS shading arm was not working.  Cycled instrument power. RTS 1542 GMT.

8/10/11

SD

RSS shadow band was dead this morning.  Cycled the instrument power to no avail at 1530 GMT.  At 2030 GMT Pat noticed that the shadow band was working again.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  At 1517 GMT found RSS computer not running.  Powered up computer.  Instrument is now active and shadowband is shading.  RTS at 1532 GMT.

8/8/11

SD

RSS Shading band was not moving. Cycled instrument power. RTS 1621 GMT.

8/4/2011

PAD

Purge Gas Generator located in Optical trailer that services RSS, SWS, and SORTI is not operating correctly.  Condensation is getting past unit and into lines going to instruments.  Unit needs to be removed and sent for repairs.  At 1520 GMT, shut down compressor and removed unit.  Whatman model FR-IR 75-45-12VDC, SN 7545-0236.  RTS at 1600 GMT.

8/1/11

PAD

Observed shadowband is not active.  Opened enclosure and found transmit light on bit driver is on solid.  Four red LEDs are lit on control/comm board.  Powercycled instrument and it will now communicate with computer.  Shadowband is now active and shading correctly.  RTS at 1431 GMT.

8/31/11

SPD

RSS shadow band was dead.  Found that the instrument power switch/circuit breaker had tripped to the off position. Turned switch to on position and it tripped again. Once more turned switch to on position and it stayed on. RSS powered up and shadow band shaded the diffuser.

8/10/11

SPD

RSS shadow band was dead this morning. Cycled the instrument at 1530 GMT to no avail.  At 2030 GMT, noticed that the RSS shadow band was working as advertised but at 2045 it was dead again.

08/08/11

SPD

RSS shading band was not moving.  Cycled instrument power. RTS 1621 GMT.

8/1/11

PAD

Observed shadowband is not active.  Opened enclosure and found transmit light on bit driver is on solid.  Four red LEDs are lit on control/comm board.  Powercycled instrument and it will now communicate with computer.  Shadowband is now active and shading correctly.  RTS at 1431 GMT.

7/26/11

CAM

Last successful SDS collection was at 0607 GMT.  Instrument is operating as it should, but computer shows INIT: Version 2.86 Reloading instead of normal login prompt.  Rebooted computer at 1830 GMT.  Computer booted up to normal login prompt.  The following collection was successful at 1907 GMT.

6/3/2011

JEM

The RSS has not been working since 5/20/11.  The mentor mailed 2 new MAX3232 chips and two microprocessor chips as well as 2 optical bit drivers.  Replaced the 4 chips on the logger board.  Reinstalled the logger board @1600 GMT.  Also went ahead and replaced both optical bit drivers.  Powered on the system.  Was able to get a few HELLO's from the RSS after powering the system on.  Notified the mentor and he was going to get in remotely and see if he could get the RSS collection going again.  RTS @1615 GMT.  At 1804 GMT the RSS was observed rotating, which it had not done for quite some time.

5/25/11

PAD

Last SDS collection was at 0209 GMT.  At 1413 GMT, nothing was shown on monitor after waking up computer with mouse.  After pressing ctrl/alt, INIT: version 2.86 reloading was shown instead of normal login prompt.  Rebooted computer and login prompt is now shown.  RTS at 1428 GMT.

5/23/11

PAD

Shadowband is not active.  Instrument bit driver transmit light is on solid and computer bit driver receiver light is solid.  Power cycled instrument.  Transmit light came on solid within seconds of instrument restart.  Last data shown for instrument on ARM home page is for 5/3/11.  Notified instrument mentor.  Troubleshooting at this time.  RTS at 1447 GMT.

4/27/11

PAD

Instrument has not produced data since 4/22/11.  Power cycled instrument and rebooted computer on 4/25/11.  Shadowband is not active.  Connected laptop directly with instrument and found no issues.  Checked comms with loopback test and found no issues.  Connected RSS computer to instrument and was able to check comms using Minicom program.  No issues were found.  Rebooted computer and power cycled instrument.  Notified instrument mentor.  RTS at 2128 GMT.

4/25/11

PAD

Last RSS data shown was on 4/22/11 at 1640 GMT.  Shadowband is not active.  Powercycled instrument and rebooted computer.  RTS at 1832 GMT.

3/7/11

PAD

RSS computer needs to be moved to another location within the Optical trailer to accommodate new instruments.  Shut down computer and relocated.  Powered up computer and verified instrument was still operating correctly.  RTS at 1932 GMT.

1/10/11

PAD

Nothing is shown on monitor.  Instrument is shading.  Rebooted computer.  Computer restarted correctly and log in prompt is now shown.  RTS at 1541 GMT.

12/30/10

PAD

Shut off enclosure AC unit at 1601 GMT.

12/1/10

CAM

Site power needs to be shut off to install the transformer and run power to the new office building.  Instruments need to be shut down gracefully.  Shut the RSS computer off at 1424 GMT.  Restarted after power was restored and RTS at 1842 GMT.

11/18/10

CAM

AC power needed to be secured to drill pier holes for new construction near the RCF trailer.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 1437 GMT.  Restarted instrument after AC power was restored.  RTS at 1746 GMT.

11/10/10

CAM

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 2156 GMT.  Restarted instrument after AC power was restored.  RTS at 2300 GMT.

9/15/10

PAD

Shadowband is not active.  Power switch for control/ comm board had tripped.  Reset switch.  RTS at 1919 GMT.

9/13/10

PAD

RSS shadowband is not active.  Last file shown is on 9/10/10.  Connected to instrument control/communication board and was able to get "Hello" response.  Tested comms between instrument and computer and found no problem.  Notified instrument mentor and was advised to wait until 30 minutes after the hour when computer sends a command to instrument to start collections.  Observed instrument shading at 1635 GMT.

8/2/10

PAD

Bit drivers provided by mentor were installed on RSS system while troubleshooting problem.  Mentor request bit driver removed from system and returned.  Shut instrument down at 1344 GMT.  R/R bit drivers.  RTS at 1359 GMT.

7/29/10

PAD

Control/comm board was removed on 7/28/10 to install new microprocessor chips.  Board needs to be installed.  Installed board and connected to laptop.  Powered up board and was able to get HELLO prompts back from board.  Connected to RSS computer and able to communicate with board using RSS computer.  Instrument soon started shading.  Contacted instrument mentor.  RTS at 1459 GMT.

7/28/10

PAD

RSS has experienced sporadic operation for several weeks.  Previous troubleshooting has been unable to isolate what is causing sporadic operation.   Mentor has sent replacement RS232 chips and microprocessor chips to install on board to see if that corrects the problem.  Chris Martin and I installed replacement RS 232 chips on the control/comm board inside instrument enclosure.  Unable to communicate with laptop or RSS computer.  Removed board and took to lab.  Original microprocessor chips that were previously removed from board were installed.  There was also a small amount of corrosion on some of the sockets microprocessor boards are seated.  Corrosion was removed and chips were installed.  Still unable to connect to control/comm board with laptop or RSS computer.  Removed board to install new microprocessor chips sent by mentor.

7/22/10

PAD

RSS has been experiencing communication issues between computer and instrument for some time.  Previous troubleshooting has failed to isolate and identify specific problem.  Mentor has sent replacement bit drivers, RS232 chips, and microprocessor chips to try and correct the problem.  Mentor requests bit drivers replaced first.  Receive light is on solid most of the time at computer, transmit light at instrument.  Shut down control/comm board on instrument.  Computer department and I tested fibers with meter and found no problem.  Tested comms with laptop and found no problem besides first letter entered was shown as wrong letter.  No failures after that.   Installed bit drivers provided by mentor.  Tested comms with only one letter shown wrong.  This only happens when receive light is on solid on bit driver.  Unable to get comms test to repeat wrong letters or symbols.  Connected RSS computer to instrument and powered up board.  At 1652 GMT, observed instrument shading correctly.  Notified instrument mentor.  Will continue to monitor.  RTS at 1652 GMT.

7/15/10

PAD

Shadowband is not operating.  Opened enclosure and shut off control/comm board.  Connected laptop to board and powered up.  Unable to connect with laptop.  Comms test good between RSS instrument and computer.  Reconnected instrument to RSS computer and powered up board.  Notified mentor.  Waiting on instructions.

7/15/10

PAD

Shadowband is not operating.    Powercycled control/comm board at 1518 GMT.  RTS at 1533 GMT.

7/14/10

PAD

Shadowband is not operating.  Opened enclosure and verified control/comm board has power.  Powercycled board and observed at 1757 shadowband was operating.  RTS at 1545 GMT.

7/13/10

PAD

RSS control/comm board was removed on 7/12/10, see CM report for details.  Board was taken to lab to have chips reseated.  Lab technician reseated chips and found some suspect solder joints on board which were repaired.  Board needs to be installed in RSS instrument enclosure to test.  Installed board and reconnected all wires.  Powered up unit and shadowband started shading.  Notified instrument mentor so data quality could be checked.  RTS at 1418 GMT.

7/12/10

PAD

At 1442 GMT, observed shadowband was not operating.  Power cycled bit drivers and instrument control/comm board at 1442 GMT.   At 1610 GMT, bit driver at instrument shows transmission light solid for ~7 seconds, flickering a couple of seconds, then repeating cycle.  Bit driver at computer mirrors pattern with receive light.  R/R bit drivers with no change.  Able to connect directly with instrument using laptop with ProComm . Instrument will respond with Hello prompts but program will freeze or display symbols.  Computer department and I bypassed comms and connected instrument directly to computer using Ethernet cable with serial adapters.  Same results were found using RSS computer to communicate with instrument as with laptop.  After consulting with mentor, removed control/comm board and took to lab to have chips reseated.  RTS at 2050 GMT.

7/9/10

PAD

On 7/8/10 at 1603 observed shadowband was not operating.  Removed cover and verified instrument control board has power.  Checked comms between computer and instrument.  Connected to control board with laptop to verify board was active.  Shut down board, reconnected comms, and powered up board.  Notified mentor who was able to remotely access computer and initialize instrument.  RTS at 1452 GMT.

6/25/10

PAD

Computer has been unable to communicate with instrument since 6/13/10.  See previous CM reports for details.  After troubleshooting on 6/25/10 and able to communicate with laptop computer, mentor is still unable to communicate with instrument using RSS computer.  Mentor request settings on bit drivers checked.  Checked switches on bit drivers.  Found computer bit driver set to DTE and instrument bit driver set to DCE, which is correct.  Power cycled instrument after consulting with mentor and who was remotely logged into RSS computer.  Computer is now communicating with instrument.  Check of instrument at 1530 GMT show instrument is now shading.  RTS at 1450 GMT.

6/25/10

PAD/CAM

Control board was installed on 6/24/10 and mentor was still unable to remotely access computer and connect to instrument.  Communication tests on comms system with two laptops indicated a faulty bit driver.  R/R bit drivers located at instrument and computer and now able to communicate without problems.   Able to connect laptop with instrument using ProComm.  After setting DTE/DCE switches, we can now communicate using laptop with instrument from Optical trailer.  RTS at 2020 GMT.

6/24/10

PAD/CAM

RSS control board was removed on 6/21/10 and sent for repairs.  Board needs to be installed.  Installed board but mentor is still unable to remotely access computer and connect with instrument.  Unable to connect to instrument with laptop computer using Hyperterminal.  Troubleshooting at this time.   RTS at 2200 GMT.

6/21/10

PAD

Instrument quit operating on 6/13/10.  Previous troubleshooting found failed bit driver and power switch shut off that supplies power to electronic boards.  Power was restored and further troubleshooting with mentor found suspected failed chips on RSS processor board.  Mentor requests board removed and sent for repairs.  Shut off power to electronic boards at 1738 and removed RSS combo/ processor board as instructed.  Board will be sent to mentor for repairs.  RTS at 1758 GMT.

6/15/10

PAD

On 6/14/10, observed RSS computer is running but shadowband is not operating.  Contacted mentor who was unable to establish contact with instrument.  At 1345 GMT, computer department and I found bad bit driver located at instrument and R/R.  Tested comms with laptop through bit drivers with no further issues found.  Later I was notified by mentor he is still unable to connect with instrument.  Tested computer serial port with no problems found.  Mentor suspects failure on RSS comms board.  Troubleshooting at this time.  RTS at 1718 GMT.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.  Powered on computer at 1356 GMT.  Shading was checked at 1440 GMT.  RTS at 1411 GMT.

5/13/10

PAD

CF lost power from 0623-0759 GMT.  Powered up RSS computer at 1336 GMT.  At 1430, checked shading and found computer was still not controlling shadowband.  at 1607 GMT, observed shadowband operating.  Unable to check shading due to sky conditions.  RTS at 1351 GMT.

1/13/10

PAD

Shadowband is not fully shading diffuser and needs adjustment.  Adjusted shadowband at 1922 GMT.  RTS at 1937 GMT.

1/12/10

PAD

Shadowband is shading diffuser but is not centered and needs adjustment.  Adjusted shadowband at 1857 GMT.  RTS at 1912 GMT.

10/5/09

CAM

Instrument mentor advised he was unable to access the RSS computer.  Status check found the computer had locked up with error: Timed out waiting for DMA.  Power cycled computer at 2155 GMT.  Notified mentor who was then able to access computer and restart instrument.  RTS at 2204 GMT.

9/16/09

PAD

Shadowband is moving too far and not shading all of diffuser pad.  Adjusted shadowband at 1756 GMT.  RTS at 1811 GMT.

8/24-28/09

Mentor

Instrument reinstalled by mentor.

2/28/08

CAM/MK

Instrument mentor requested the RSS be removed and returned for refurbishment (ECO-00661).  Performed calibration on the instrument from 1856 to 1925 GMT.  Disconnected power and communication lines and removed the entire box from the stand.  Delivered instrument to S&R to be crated and shipped next week.

12/24/07

MK

Due to a power outage the UPS has completely drained the internal batteries and is unable to restart. Reset UPS at 1430 GMT, all instruments in the Optical facility have been restarted and are currently running.

3/20/07

PAD

Storms at CF tripped UPS located in Optical trailer.  Last successful collection was at 0300 GMT.  Following instruments were affected:  AERI01, MWR, RSS, SWS, TSI-880, SORTI01, and VCEIL25K.  Facilities reset UPS at 1600 GMT.  Restarted computers that did not restart after power was restored.  RTS at 1635 GMT.

7/6/06

PAD

Performed RSS Port Cal starting at 1838 GMT.  RTS at 1901 GMT.

12/20/05

PAD

Conduit protecting fiber wires does not have the right size connector to connect to instrument shelter.  Craig Webb, Mark Klassen and I shut down instrument at 1914 GMT.  Removed fiber wires and drilled hole in instrument shelter to accommodate 1" conduit.  Attached conduit to shelter and reconnected fiber optic wires.  Instrument restarted correctly.  RTS at 1934 GMT.

12/8/05

MK

Informed by John Schlemmer that the RSS is not running. The linux-side software is complaining that the instrument isn't talking.  As the band has stopped spinning, he speculates that the power to the instrument has been interrupted (if this was just a case of a F/O failure the band would keep on spinning). After determining that the correct power was being supplied to the RSS. It appears that no power is being supplied to the combo board.  The 10 led light bar and red led never light up. I opened up the enclosure that houses the DC power source, we checked four fuses inside, all checked good.   Checked both bit drivers for proper supply voltage, power cycled both optical bit drivers. Received operating manual from Yankee Environmental Systems on 12/9/05 and it was determined that system will not run if temperature is below 0°C. Verified that the internal heater was coming on. At approximately 2000 GMT, RSS restarted and is shading correctly and data is being collected once daily.

11/21/05

PAD

SDS last collection was at 1102 GMT.  Instrument was operating and diffuser was shading correctly.  Computer shows RSS log in prompt.  SDS collection log shows no files to get with extension gz.  Notified computer department and was advised to reboot computer.  Rebooted computer at 1539 GMT.  SDS did not collect following hour.  Contacted instrument mentor and was informed instrument only collects once/ day.  RTS at 1554 GMT.

9/23/05

MR

The instrument mentor wants the port cal and LICOR calibration run on the instrument.  Performed the calibration.  RTS at 1800 GMT.

9/22/05

MR

The instrument mentor wants two calibrations ran today on the RSS to help pinpoint a problem.  Performed the first calibration at 1445 GMT.  RTS at 1530 GMT.

9/21/05

MR

The instrument mentor wants three calibrations ran today on the RSS to help pinpoint a problem.  Performed the third calibration at 2045 GMT.  RTS at 2130 GMT.

9/21/05

MR

The instrument mentor wants three calibrations ran today on the RSS to help pinpoint a problem.  Performed the second calibration at 1730 GMT.  RTS at 1815 GMT.

9/21/05

MR

The instrument mentor wants three calibrations ran today on the RSS to help pinpoint a problem.  Performed the first calibration at 1445 GMT.  RTS at 1530 GMT.

8/5/05

MR

The instrument mentor is on site and wants a calibration performed. Performed the calibration procedure.  During the procedure the calibration was suspected of being bad.  The mentor stopped the calibration and wants another one ran next week.  RTS at 1900 GMT.

5/4/05

MR

The instrument mentor reports that the portable calibrations are not working properly and the files have been corrupt for the last 3 calibrations.  He wants a special calibration ran while he logs on and watches the data.  Performed 3 calibrations with the portable calibrator while talking to the mentor on the phone.  It was determined that the calibrator is defective and he wants it shipped back to him for inspection and repairs.  RTS RSS at 1930 GMT.

4/15/05

MR

The biweekly calibration is due.  Performed the biweekly calibration.  RTS at 2000 GMT.

2/25/05

MR

The instrument mentor wants the RSS analog card adjusted as per his instructions.  Contacted the mentor.  While he was on the phone, he had me turned the analog adjustment pot 1/2 turn clockwise while he monitored the results.  He then had me perform a calibration using the portable calibrator.  RTS at 1745 GMT.

2/17/05

MR

The RSS calibration is due.  Performed the RSS calibration.   RTS at 2130 GMT.

2/3/05

MR

The RSS calibration is due.  Performed the RSS calibration.  Notified the mentor.  RTS at 2200 GMT.

12/23/04

MR

The RSS calibration is due.  Performed the RSS calibration as per mentors instructions.  RTS at 2145.

11/18/04

MR

The RSS calibration is due.  Performed the RSS calibration between 2010 GMT and 2100 GMT.  RTS at 2055 GMT.

10/14/04

MR

The instrument mentor has notified us that the instrument calibrator has not been working as expected and wants me to check the calibration bulb.  T/S found the bulb was not working.  Further T/S found that the fuse was blown.  R/R the fuse at 2015 GMT.  Performed instrument calibration at 2020 GMT with expected results.  RTS at 2045 GMT.

9/28/04

MR

The instrument mentor wants me to check and make sure the calibration control will work with the new board.  Plugged in the calibration control and the band stopped moving as expected.  Unplugged the controller and RTS the instrument at 2045 GMT.

9/24/04

MR

The instrument mentor notified me and gave me instructions about how to rewire the plug and wants the plug rewired and the instrument started.  Rewired the plug as per mentors instructions.  Powered up the instrument at 1450 GMT.  At the top of the hour the instrument started and the shadow was correct.  Will check on the instrument at solar noon to make sure the band motor does not need adjustment.  Notified the mentor.  RTS at 1510 GMT.

9/23/04

MR

The instrument mentor has sent a replacement mother board for us to change out so that the shadowing problem can be solved.  Changed out the mother board.  No visible serial numbers.  One of the plugs was not able to be plugged into the board due to the pins being different and the keys in the wrong place.  Will notify the mentor and leave the instrument off until further instruction has been received.  Notified the instrument mentor and left instrument powered down at 1530 GMT

9/16/04

MR

The instrument mentor wants us to replace the shading motor to see if this will repair the shading problem.  R/R the motor at 2015 GMT.  The instrument was shading correctly BEFORE the motor was replaced since it was past solar noon.  Will have to wait and see if the instrument continues to shade correctly in the morning. 

9/16/04

MR

The instrument mentor called and wants the motor driver and motor latch IC changed on the board in order to see if this will correct the shading problem with the instrument.  Powered the instrument off at 1718 GMT.  R/R the motor driver and motor latch IC's on the data logger.  Powered the instrument back up at 1733 GMT.  This did not cure the problem.  RTS at 1733 GMT.  T/S continues.

9/17/04

PAD

RSS band has been sporadic on shading diffuser for some time.  Instrument mentor has been troubleshooting problem with John Schmelzer who is on site.  Mentor requested processor chip changed.  Opened enclosure and changed left hand processor.  Band would still not shade diffuser.  Removed and replaced right hand processor chip.  Band will still not shade correctly.  RTS at 1529 GMT.

9/9/04

MR

The instrument mentor wants the data logger board lights to be checked and the status reported back to him.  Checked the board error lights and called the mentor.  He asked that the instrument be power cycled.  Power cycled the instrument at 1945 GMT and called the mentor.  The mentor restarted the instrument and had me check the shading.  The shading is still off.  The mentor is going to check and see what may be causing this and will contact us with further instructions.  RTS at 2030 GMT.

9/3/04

MR

Morning rounds found the instrument was not running.  The mentor contacted us and said that the collections from the instrument have been down since 9/1/04 at 1433 GMT.  T/S found that the fiber from the optical trailer to the junction box was broke.  Moved the fiber from pairs 7/8 to pairs 1/2.  Communications with the instrument where restored and the instrument was restarted by the mentor at 1830 GMT.  RTS at 1830 GMT.

9/2/04

MR

The calibration procedure is scheduled to be performed today.  Calibrated the instrument with the LICOR and the portable calibrator.  RTS at 1445 GMT.

8/9/04

MR

The instrument mentor wants the portable calibration and the LICOR calibration performed.  Performed both calibrations as per mentor’s instructions.  Notified the mentor.  RTS at 1545 GMT.

7/28/04

MR

The site power will be going down and this instrument needs to be shut down. Shut the instrument down between 1432 and 1558 GMT.  RTS at 1558 GMT.

6/30/04

MR

The power to the site was lost at 1330 GMT.  While power was out, the system was physically turned off so a controlled power up could be performed.  Power was restored to the site at 1455.  Powered up the system at 1541 GMT.  System came up and rebooted ok.  RTS at 1556 GMT.

5/27/04

MR

The RSS portable calibration is due.  Performed the RSS portable calibration.  RTS at 1825.

5/14/05

MR

The RSS is due for calibration.  Performed all 3 calibrations between 1445 and 1635 GMT.  RTS at 1645. 

4/20/04

MR

The computer was found to be off due to the power outage overnight.  Turned the computer back on at 1416 GMT.  RTS at 1431 GMT.

3/19/04

PAD

Mentor request RSS calibrated.  Started calibration at 1651 GMT and performed calibration with portable, Li-Cor, and Oriel Spectral.  RTS at 1751 GMT.

01/02/04

PAD

Mentor requested RSS be calibrated on weekly basis.  Started portable calibration procedure at 1942 GMT.  RTS at 2005 GMT.

12/26/03

PAD

The RSS is due for the portable calibration today.  Started calibration procedure at 1923 GMT.  RTS at 1948 GMT.  Instrument operated normally after calibration.

12/19/03

MR

All 3 calibrations need to be run on the instrument.  Performed all 3 calibration procedures on the RSS.  RTS at 2200 GMT.

12/2/03

MR

The instrument mentor wants this instrument power cycled and the internal battery unplugged during the power cycle.  Powered down the instrument at 1622 GMT.  Unplugged the internal battery and allowed the instrument to set for 10 min.  Plugged the internal battery back in and powered up the instrument.  RTS at 1637 GMT.

11/21/03

MR

The RSS is due for the portable calibration today.  Performed the portable calibration procedure.  RTS at 1530 GMT.

11/6/03

MR

The RSS calibrations are due.  Performed all 3 calibration procedures.  RTS at 1660 GMT.

8/1/03

JEM

Calibration needs to be performed using the portable calibrator.  Performed the scheduled calibration on the RSS from 1535 GMT to 1555 GMT.  The RSS was returned to service and was operating normally after the calibration was performed.  RTS @1555 GMT.

6/26/03

JEM

The RSS needs to be calibrated using the portable calibrator.  Disabled and calibrated the RSS from 1445 GMT to 1510 GMT.  Removed the calibrator and returned the RSS to service @1514 GMT.  RTS @1515 GMT.

6/2/03

JEM

The RSS shadowband is not rotating.  Powered off the RSS for approximately 20 minutes.  Powered the RSS back on @1501 GMT.  Notified the mentor so the RSS could be remotely initialized.  RTS @1510 GMT.

5/27/03

JEM

RSS calibration needs to be performed.  Calibrated the RSS using the Portable Calibrator from 1815 GMT to 1900 GMT.  RTS @1900 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

1/27/03

MR

The instrument mentor has requested that the computer be checked due to it has lost communications.  T/S found the computer was turned off.  Powered on the computer.  RTS at 1815 GMT.

7/26/02

CT

The Shadow band was continuously trying to rotate. It was hitting the west side of the sensor base. The mentor was contacted by email. The mentor requested that the instrument be power cycled. The instrument was power cycled at 1630 GMT. RTS at 1645 GMT.

7/1/02

CAM

A fiber optic cable between the RSS computer and instrument was discovered on 6/24.  The fiber had been chewed through by mice.  The fiber was replaced on 6/25 and the instrument needed to be restarted.  Notified instrument mentor and the RSS was restarted at approximately 1615 GMT on 7/1.

6/18/02

MR

The instrument mentor has had problems with communications with this instrument since the storm on 6/13/02.  T/S found the external GFCI breaker was tripped.  Reset the GFCI and I was able to see the data logger board start to flash the bit driver transmit light.  Notified the mentor.  Awaiting replay from mentor.

6/14/02

MR

Power restored.  Powered on computer and notified mentor the instrument needed to be restarted.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/7/02

KT

The instrument is down.  The shadow band is resting on the West side. The instrument mentor was contacted by e-mail about the problem. Troubleshooting is in progress. RTS at 1412 GMT.

5/29/02

CT

The sensor was not being shaded properly. The mentor requested at solar noon that the vertical bracket be checked for shading . There was not shading by the vertical bracket. Mentor requested that the shadow band be adjusted after this step was taken.  The shadow band was adjusted at 1920 GMT. RTS at 1935 GMT.

5/21/02

CT

The RSS instrument is down.  Was notified by Chris Martin that the RSS mentor requested that the IBM Aptiva computers be powered on. At 1415 GMT the RSS computers located in the Optical trailer were powered on. RTS at 1430 GMT.

5/15/02

CAM

Instrument mentor advised that the computer which controls the U1000 spectrometer is not responding to pings.  T/S found that the computer was powered off.  Powered the computer on at 1822 GMT and verified that it could now be pinged over the network.  Notified mentor.

8/20/01

MR

The power was lost to the site on 8/15/01 from 0030 GMT to 0400 GMT.  This instrument needs to be rebooted and restarted. Powered up the computer at 1428 GMT.  Tried to call Lee Harris and have him reboot the instrument but was unable to reach him at this time.  Will keep trying until he is reached.

6/21/01

Mentor

Instrument mentor (Jerry Berndt) on site all this week to reinstall the RSS.  Installation/testing completed and system RTS at approximately 2200 GMT on 6/21.

9/18/00

MR

The mentor has removed the instrument. No further PM at this time.

9/11/00

KT

At 1509 GMT called mentor and left a voice mail message stating that the RSS instrument was power cycled on 9/8/00 per request and the shadow band was not rotating. No further action taken, awaiting mentor reply. RTS at 1523 GMT.

9/8/00

KT

At 1840 GMT mentor requested the reboot of the instrument. At 1841 GMT rebooted the instrument. RTS at 1855 GMT.

8/17/00

KT

Mentor stated that they could not get any data from the RSS computer. At 1600 GMT found that the RSS computer was locked up and that there were no transmit or receive lights being lit on the bit driver. At 1604 GMT reseated the power plug on the bit driver. At 1605 GMT rebooted the computer. At 1607 GMT the computer came on line with a RSS-CF login prompt. The transmit and receive lights were also flashing on the bit driver. RTS at 1615 GMT.

7/28/00

CAM

Jim Schlemmer reports that he has not been able to communicate with the RSS instrument since 7/26. T/S found that the DC power connector on the fiber optic bit driver was not fully seated. Fully seated the power connector at 1925 GMT. Jim verified that communications were restored at 1927 GMT. RTS status O.K.

6/1/00

ZW

For construction reasons the RSS was powered down, however it was not powered back up again. Powered the RSS computer back up at 1636 GMT. Status OK. RTS 1650 GMT.

3/27/00

MR

Jess Smith notified me that the RSS was not communicating with the SDS. Rebooted the SPARC station. Connection was restored.

3/10/2000

CAM

Jim Schlemmer notified Site Ops that the RSS SparcStation locked up while he was trying to restart the RSS remotely. T/S found that the Sparc was locked up with a blank screen. Power cycled the Sparc at 2125 GMT. Notified Jim that the system was now ready to be reset. Status check at 2150 GMT found that the system was fully operational. RTS status O.K.

3/10/2000

CAM

The RSS shadow band was not rotating. The band was stopped hard at the midway point. Power cycled the instrument at 1750 GMT. The system must be restarted remotely by the mentor. Left voice and email with the mentor to restart. System remained off line as of 2100 GMT. Waiting for restart.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will need to be shut down. Shutdown RSS at 1435 GMT, restarted RSS at 1700 GMT. RTS at 1700 GMT.

11/24/99

CAM

System Administrator notified Site Ops that the RSS Sun computer was not visible on the network. The instrument mentor had also indicated he was not able to access the computer. Logged on to the Sun computer and everything appeared normal. The Ethernet transceiver link light was lit and the transmit and receive lights were flashing. Reset the Ethernet port at 1745 GMT. The Sun computer communications with the network were then restored. RTS at 1800 GMT.

10/26/99

TMG

Jim Schlemmer reports that he can not communicate with the RSS. T/S of system indicates that the SUN computer port "A" is not working and no signal is being sent to the monitor. Technician rebooted computer at 1615 GMT as per Jim Schlemmer. System RTS at 1638 GMT. Jim Schlemmer checked status of RSS. Status O.K.

10/15/99

KT

At 1858 GMT the outside compressor was powered off due to leaking connecting putty from an elbow connection on the side of the compressor.

7/19/99

MW

The RSS was installed today. Hooked power and communications up 1600-1620 GMT. The instrument mentor is on site verifying proper operation at this time.

4/7/99

CAM

RSS instruments are removed from the site at this time for calibration.

11/23/98

MW

The instrument mentors were on site today to remove both RSS systems from service for troubleshooting and maintenance. Shut both systems down at 1500 GMT and disconnected power and communications.

11/02/98

SS

Neither RSS1 or RSS2 shadow bands are not rotating. No CM performed at this time due to flooding conditions.

10/13/98

MR

The shadow band is not rotating. Called the mentor and he reported that the unit is currently out of service and will be removed.

10/9/98

ZW

The Shadow band is not rotating at this time the Mentor has stopped it because of a problem with the shutter.

10/6/98

CAM

Instrument mentor indicates that the shutter appears to be stuck in the open position beginning 9/5/98 at 2101 GMT. Mentor requested that the shutter power plug be temporarily disconnected. Removing power from the shutter should cause it to shut immediately unless there is a physical problem with the shutter. Opened the optics housing and disconnected the shutter plug at 1817 GMT. Phoned mentor who advised that from viewing the data the shutter appears to remain open. Reconnected the shutter plug, closed the optics housing and RTS at 1832 GMT. The shadow band stopped rotating when the mentor connected remotely and due to an unrelated problem he is unable to restart the band at this time. Mentor will contact Site Ops at a later time with further instructions.

10/2/98

MR

Power cycled the data logger at 1520 due to band not running.

10/1/98

CAM

Instrument mentor requested that Site Ops power cycle RSS 103. It appeared as though the shutter may have been stuck open and the system hung up while mentor was logged on remotely. Power cycled RSS master control switch at 1957 GMT. Mentor restarted remotely and RTS at 2005 GMT.

9/30/98

CAM

Instrument mentor notified Site Ops that changing the cooler on 9/28/98 solved the temperature problem. Mentor requested Site Ops return the old cooler to him. Returned Peltier Cooler SN: R-75364, Model: AAC-141-4XTHD, Mfg: Electrografics Intn'l Corp.

 9/28/98

CM/GS

Instrument mentor advised that the RSS Peltier Cooler had failed. Mentor shipped a new cooler to replace the failed unit. Notified mentor that the cooler was received and mentor stopped the RSS data acquisition at 1910 GMT. Powered the RSS master control switch off at 1921 GMT. Opened the optics housing and removed the old *Peltier Cooler at 1930 GMT. Installed replacement cooler and set thermostat control knob to 20. Resealed optics housing at 2010 GMT. Powered the RSS master control switch on at 2015 GMT. Notified mentor that the swap was complete and RSS was ready to be restarted at 2020 GMT. Mentor will monitor the temperature to verify the problem is solved. *Old - R-75364 *New - N/A

9/15/98

CAM

RSS mentor advised that the RSS Peltier cooler on RSS #103 is not operating properly. He further requested that the cooler be removed from AC power for a couple of hours to see if it may be thermally locked up. Unplugged the cooler from AC power at 1511 GMT. Plugged cooler back in to AC power at 1810 GMT.

7/22/98

MW

The RSS shadow band was not rotating, probably due to the power bump the CF experienced this morning. Called the instrument mentor at 1515 GMT and informed him that the RSS needed to be restarted. The band was rotating as of 1535 GMT.

7/7/98

MW

Found the chiller without power. The chiller AC power cord was plugged into the lower outlet. Plugged the cord into the upper outlet and the chiller restarted at 1830 GMT.

7/7/98

MW

The RSS shadow band is not rotating. Called the instrument mentors and left a message to restart the RSS at 1835 GMT.

6/26/98

MW

The RSS instrument mentor requested the RSS be calibrated this week. Performed RSS calibration from 1335 GMT to 1410 GMT. RTS at 1415 GMT.

5/22/98

MW

Both RSS desiccant tube (internal+external) desiccants were pink. Replaced the external desiccant tube at 1755 GMT. Will monitor the internal desiccant to see if it changes back to blue.

5/13/98

MW

The instrument mentor was on site today to perform system calibration procedures. The RSS was stopped at 1325 GMT and then restarted at 1403 GMT after calibration procedures were completed.

5/12/98

MW

The RSS mentor was on site today to repair a broken cooling unit and perform system calibration procedures. The RSS was stopped at 1915 GMT and a new cooling unit was installed. Then calibration procedures were performed until 2110 GMT when the system was restarted.

4/9/98

CM/DN

Instrument mentor requested Site Ops assistance in calibrating the RSS. Mentor disabled the RSS at 1750 GMT. Site Ops set up the HgCd lamp and powered on at 1800 GMT. Mentor completed the wavelength calibration at 1810 GMT. Lowered the optics housing in preparation for absolute calibration at 1820 GMT. Positioned the Licor Calibrator over the fore-optic assembly and notified mentor for dark count calibration. Mentor completed dark count calibration at 1832 GMT. Site Ops powered the Licor Calibrator on at 1833 GMT. Mentor completed the absolute calibration and Site Ops powered the calibrator off at 1848 GMT. Site Ops raised the optics housing and leveled the sensor at 1910 GMT. Mentor restarted the RSS and RTS at 1918 GMT.

4/2/98

MW

The RSS shadow band was not rotating, possibly caused by a power bump since the system is not on UPS power. Power cycled the instrument at 1540 GMT. The instrument mentor restarted the system remotely at 1558 GMT. RTS.

3/20/98

SS

Work order 2741. The mentor indicated that he would like the shading checked when the weather permits. Checked the shading in the A.M. during daily checks and it was shading properly. Checked the shading again at 1940 and the west side of the sensor was slightly exposed to direct sunlight. Adjusted the shadow band for full coverage of the sensor at 1940 GMT. RTS at 1940 GMT.

3/20/98

SS

The mentor requested that the shadow band alignment be checked when weather permits after instrument installation. Checked shadow band alignment. Found it to be aligned correctly.

3/17/98

MW/SS

The instrument mentor did not install the outer shield on the RSS optics enclosure due to wet conditions at the time the instrument was installed. Installed outer shielding on the RSS optics enclosure 1915-1945 GMT.

3/16/98

MW

The instrument mentor installed the RSS today. RSS operational as of 1820 GMT.

2/19/98

MW

The instrument mentor wanted to perform some calibration checks and to remove the RSS from service for upgrading. Stopped data collection at 1550 GMT. Performed mercury cadmium lamp calibration check 1555-1605 GMT. Performed absolute calibration with the optical radiation calibrator 1627-1645 GMT. Powered the RSS off at 1648 GMT and removed it from service.

1/23/98

MW

The external desiccant tube desiccant was pink. R/R desiccant tube at 1530 GMT.

1/16/98

KT

Performed weekly at 1740 GMT.

1/9/98

MW

The RSS shadow band was not rotating properly. The band was vibrating back and forth in the home position. Powered the RSS off at 1530 GMT. Observed no water inside the motor connectors. Blew air into connectors to ensure that there was not any water inside which could not be visually detected. Powered the RSS back on at 1540 GMT. Called the instrument mentor to have the RSS initialized. Mentor initialized the RSS at approximately 1545 GMT. The shadow band began vibrating once again in the home position. Powered the RSS off at 1550 GMT. R/R the shadow band motor. Powered back on at 1555 GMT. The instrument mentor initialized the RSS at approximately 1558 GMT. The band began rotating normally after the motor exchange. RTS at 1600 GMT. Could not fine adjust the shadow band for proper shading due to overcast conditions. Need to fine adjust next time weather conditions permit.

1/6/98

KT

The RSS shadow band was stuck on the west side of the motor drive assembly and was just vibrating without rotating. At 1645 GMT the RSS was powered off and the motor cable was removed. There was no water observed in the connector. Used compressed air to blow out the connector and reinstalled. Restored power at 1648 GMT. At 1651 GMT the mentor was called and notified of action taken and was asked to remotely reset the instrument. At 1825 GMT the mentor phoned to advise the reset had not worked and requested Site Ops spray the motor cable with WD40 and dry again. Powered system back off at 1830 GMT. Removed cable and sprayed with WD40. Allowed cable to dry and reinstalled. Mentor attempted to reset but once again the shadow arm would only vibrate. Weather conditions were drizzly so the system was left in present state. T/S will continue when the weather clears as per mentor.

12/29/97

MW

The RSS shadow band was not rotating. Instead, it was vibrating in the home position. Removed the motor cable from the motor at 1510 GMT. Observed a small amount of water inside the connector. Left the cable off to air dry. Powered the RSS off at 1519 GMT and reconnected the motor cable. Powered back on at 1520 GMT. Called the instrument mentor and left a message to have him start the RSS.

12/15/97

MW

The RSS instrument mentor called and reported a problem with the RSS which started on 12/13/97. Found the shadow band was vibrating back and forth against the MFR sensor housing. Power cycled the RSS at 1500 GMT as per mentor instruction. Mentor started the RSS remotely but the shadow band would not go to the home position. Powered the RSS off at 1535 GMT and placed the shadow band in the vertical position and powered back on as per mentor instruction. This time the band began vibrating rapidly back and forth in the vertical position. Powered the RSS off at 1600 GMT and disconnected the motor cable from the motor. Observed water inside the connectors. Blew the connectors out with canned air and wiped up excess moisture with a towelette. Powered the RSS back on at 1615 GMT. The mentor restarted the system at 1630 GMT and reported everything looks OK on his end. The shadow band is now rotating normally. RTS at 1635 GMT.

11/25/97

CAM

RSS shadow band was not rotating during daily PM. The UPS in the optical trailer indicated there had been a power bump at 2034 GMT on 11/24. Phoned RSS contact Joe Michalsky at 1525 GMT on 11/25 and advised of the situation. Joe advised he would restart the system.

10/31/97

CAM

Performed the weekly PM. Status O.K.

10/21/97

MW

The CF experienced a power bump this morning at approximately 1335 GMT. This reset the RSS electronics and the shadow band stopped rotating. Called the RSS mentor at 1440 GMT and informed him that the instrument was reset and needed to be restarted. Mentor restarted the RSS remotely at 1515 GMT.

10/17/97

MW

The RSS external desiccant tube desiccant was pink. R/R external desiccant tube at 1440 GMT.

10/10/97

CAM

Performed the weekly PM checks. Status O.K.

9/24/97

MW

The RSS instrument mentor requested that the CCD controller board CPU (HD63C03YP) and the shadow band controller board CPU (HD63C03YP) be changed out due to a communications problem between the two controllers. Called the RSS instrument mentor at 1530 GMT and he turned off the RSS data acquisition process at 1534 GMT. Powered the RSS master switch off at 1535 GMT. R/R both CPU's and powered the master switch back on at 1553 GMT. Called the RSS mentor at 1555 GMT and informed him that the procedure had been completed so data acquisition could be resumed.

9/8/97

MW

The RSS instrument mentor called site operations and informed ops about a possible cooler problem with the RSS. Found the external fan on the cooling unit was not working. Checked the AC outlet, the cooling unit is receiving power. Switched the power cord from the upper outlet to the lower outlet with no effect. Powered the RSS off at 1504 GMT. Checked all fuses. Fuses ok. Powered back on at 1515 GMT. Opened the West panel and noticed the internal fan is working but the external fan is not working (separate power supplies). Informed instrument mentor of the situation and he suggested the RSS be powered off until he can service the instrument. Suspect one of the power supplies to the cooling unit is bad. Powered the system off at 1600 GMT.

9/3/97

MDW

Due to a Kay Electric transformer going down in Billings on 9/2, the site lost all power at around 2030 GMT. The extended power outage caused the optical trailer UPS to drain its batteries and all instruments lost power around 2130 GMT on 9/2. Instrument check on 9/3 after power was restored found that the shadow band was not rotating. Phoned instrument mentor who advised to power cycle the instrument. Power cycled RSS at 1502 GMT on 9/3. Instrument mentor restarted the RSS remotely at 1506 GMT. System RTS status O.K. but unable to verify proper shading due to ovc conditions.

8/29/97

MW

The flow meter inside the RSS electronics enclosure was reading slightly low (0.8). Adjusted the flow meter valve at 1435 GMT. The flow meter now reads 2.5. The pressure meter also came up to 4.9. Both readings are within specifications.

8/29/97

MW

The desiccant inside the RSS desiccant tube was pink. R/R the desiccant tube at 1432 GMT.

8/26/97

MW

The RSS shadow band stopped rotating due to a power bump at 1512 GMT. Called the instrument mentor at 1530 GMT and informed him of the situation. The RSS was re-started remotely at 1549 GMT. RTS.

8/25/97

MW

The RSS shadow band is not rotating. Called all four instrument contacts at 1500 GMT and notified them of the situation.